# Green Edge AI: A Contemporary Survey

### v1

@@ Abstract
Artificial intelligence (AI) technologies have emerged as pivotal enablers across a multitude of industries, including consumer electronics, healthcare, and manufacturing, largely due to their significant resurgence over the past decade. The transformative power of AI is primarily derived from the utilization of deep neural networks (DNNs), which require extensive data for training and substantial computational resources for processing. Consequently, DNN models are typically trained and deployed on resource-rich cloud servers. However, due to potential latency issues associated with cloud communications, deep learning (DL) workflows are increasingly being transitioned to wireless edge networks in proximity to end-user devices (EUDs). This shift is designed to support latency-sensitive applications and has given rise to a new paradigm of edge AI, which will play a critical role in upcoming sixth-generation (6G) networks to support ubiquitous AI applications. Despite its considerable potential, edge AI faces substantial challenges, mostly due to the dichotomy between the resource limitations of wireless edge networks and the resource-intensive nature of DL. Specifically, the acquisition of large-scale data, as well as the training and inference processes of DNNs, can rapidly deplete the battery energy of EUDs. This necessitates an energy-conscious approach to edge AI to ensure both optimal and sustainable performance. In this paper, we present a contemporary survey on green edge AI. We commence by analyzing the principal energy consumption components of edge AI systems to identify the fundamental design principles of green edge AI. Guided by these principles, we then explore energy-efficient design methodologies for the three critical tasks in edge AI systems, including training data acquisition, edge training, and edge inference. Finally, we underscore potential future research directions to further enhance the energy efficiency of edge AI.
@@ I-A Background
Six decades after its inception at the 1956 Dartmouth Conference [1], artificial intelligence (AI) demonstrated its potential to surpass human intelligence in 2016. This was evidenced when a computer program named AlphaGo [2] triumphed over a top-ranked human Go player for the first time, which not only signified a landmark moment in AI history but also catalyzed a resurgence in AI interest and development. The AI renaissance has also been fueled by advancements in computing hardware, the proliferation of big data, and the advent of deep neural networks (DNNs). As a result, AI technologies have made remarkable breakthroughs across a multitude of disciplines, including but not limited to, computer vision [3], natural language processing (NLP) [4], healthcare [5], manufacturing [6], and financial technology (FinTech) [7]. Consequently, AI is rapidly emerging as a key engine of future productivity and has been recognized as a national strategic priority by major global economies, including China, the United States, and the European Union.
@@
The success of AI is underpinned by the power of advanced data analytics, which typically entails substantial computation overhead [8]. Traditionally, AI computation workloads initiated by end-user devices (EUDs) are offloaded to the Cloud, which boasts virtually unlimited computational resources [9]. However, the prevalent cloud AI paradigm falls short in supporting latency-critical applications. This is due to the multi-hop routing and network congestion between EUDs and remote public cloud servers, which can result in round-trip time (RTT) of tens to hundreds of milliseconds [10]. Furthermore, a cloud server can become inundated by a high volume of concurrent AI service requests, leading to scalability issues [11]. Concentrating all computational resources in a few cloud data centers also exposes the system to single-point failures and security threats. Moreover, data collected from EUDs may be at risk of privacy breaches [12]. These significant limitations of cloud AI have necessitated the exploration of alternative solutions for delivering AI services.
@@
Gratefully, the drawbacks of cloud AI can be complemented by the emerging mobile edge computing (MEC) technologies, which integrate cloud-like functionalities in the radio access network (RAN) near EUDs [13]. The synergy between MEC and AI enables a major paradigm shift, from cloud AI toward edge AI [14, 15], where the AI computations of EUDs can be offloaded to a wirelessly connected MEC server. As such, the data communication latency can be trimmed significantly to the millisecond level [16], which is essential for real-time interactive AI applications such as mixed reality, smart robots, and autonomous vehicles. Edge AI was identified by Huawei, a leading provider of telecom infrastructure and devices, as one of the six pillars in the sixth-generation (6G) wireless networks [11]. The International Telecommunication Unition (ITU-R) has also defined integrated AI and communication as a key usage scenario of IMT-2030 (6G) in June 2023 [17]. Since 6G is expected to rollout by 2030, edge AI technologies are foreseen to receive unprecedented attention in the near future as predicted by the Gartner Hype Cycle [18]. Apart from latency reduction, edge AI also enjoys many other benefits over cloud AI [19, 20]. On the one hand, the relatively low investment of MEC infrastructures allows their deployment at scale so that the massive AI service demands in 6G can be accomodated efficiently. On the other hand, edge AI inherits the privacy and security enhancements in MEC, since the computing infrastructures are highly distributed and can be privately owned. For instance, without relying on the cloud service providers, user data can be kept and processed inside the local domains, such as university and enterprise networks, reducing the chance of information leakage. Nevertheless, despite all these merits, there exists a significant energy challenge of edge AI, which is caused by the rapidly growing power consumption of AI workloads and the limited energy resources over the wireless edge network [21].
@@ I-B The Energy Challenge in Edge AI
As complex DL processing tasks move away from the Cloud [32], edge AI is emerging as a more sustainable solution for intelligent mobile applications. This is largely due to the use of low-power edge computing devices, which saves the high energy costs associated with operating cloud data centers. However, the energy challenge remains a significant issue in edge AI, and in fact, becomes even more pronounced [19]. Firstly, the number of EUDs requiring AI services, such as smartphones, personal computers, and vehicles, is increasing at an astonishing rate. A market analysis report by Tractica predicted that the annual worldwide shipment of AI-enabled EUDs will reach 2.6 billion units by 2025, representing a 16-fold increase compared to 2018 [33]. The widespread adoption of intelligent mobile applications will undoubtedly exacerbate the energy crisis of edge AI. Secondly, EUDs are primarily powered by batteries, the development of which significantly lags behind AI algorithmic innovations. Consequently, if compute-intensive DL algorithms are executed on EUDs, they can quickly deplete the battery energy. This not only hinders the provision of stable AI services but may also disrupt other regular system operations. While this issue can be mitigated with MEC infrastructures, it is challenging to provide sufficient resources to meet the heavy AI computation workloads offloaded from a large number of EUDs. Thirdly, the collaboration between EUDs and MEC servers involves frequent two-way wireless communications [34]. As a result, the energy consumption of communication contributes significantly to the overall energy consumption of edge AI systems, in addition to the computation energy consumption in the mobile AI paradigm, where EUDs handle all the computation workloads.
To provide high-quality, low-latency, and sustainable AI services to EUDs, it is crucial to develop dedicated solutions to address the energy challenge in edge AI. This forms the primary motivation for this survey on green edge AI. As the first survey of its kind, our focus will be on reducing energy consumption at EUDs, given that MEC infrastructures are likely to have stable power supplies.
I-C Related Survey Papers and Our Contributions
@@
Ref.
AI Tasks
Energy Efficiency- oriented
Device- edge Cooperation
Overview
Dataset Acqui.
Model Training
DNN Inference
@@
[42] $\bm{\times}$ $\bm{\checkmark}$ $\bm{\checkmark}$$\bm{\times}$ $\bm{\checkmark}$ Emerging trends, tools, and techniques for efficient DNN inference on EUDs, including designs of architectures, algorithms, hardware, and accelorators
[43] $\bm{\times}$ $\bm{\checkmark}$ $\bm{\checkmark}$ $\bm{\checkmark}$$\bm{\times}$ A systemaic overview on green deep learning technqieus, including energy- efficient training, energy-efficient inference, and efficient data usage strategies
@@
Ours $\bm{\checkmark}$ $\bm{\checkmark}$ $\bm{\checkmark}$ $\bm{\checkmark}$ $\bm{\checkmark}$ Comprehensive literature review on EE-oriented design approaches for edge AI, including training data acquisition, distributed edge training and edge inference
There are several recent survey papers on edge AI [14, 15, 16, 34, 35, 36, 37]. Specifically, a survey on edge AI was first presented in [14], which provides an overview of the key architectures and technologies for training and inference at the mobile edge network. Applications and open challenges of the synergy between DL and MEC were discussed in [16]. The communication challenges of edge AI were analyzed in [34], for which, numerous communication-efficient techniques were introduced. Similarly, practical edge AI techniques and their interplay with advanced wireless communication systems were elaborated in [35]. The challenges of edge AI were also examined from the general perspective of resource scarcity in [36]. Scalable and trustworthy edge AI systems, built upon the latest innovations in 6G and distributed machine learning (ML) technologies, were discussed in [15]. Additionally, the authors of [37] identified the fundamental components of edge AI as edge caching, edge training, edge inference, and edge offloading, and summarized the respective research results.
@@
This paper presents a timely and comprehensive literature review on energy efficiency (EE)-oriented design approaches for edge AI. The novel contributions of this study, compared with existing survey papers on edge AI, are summarized in Table I. Specifically, we first introduce the three key tasks of edge AI, namely 1) data acquisition for centralized edge learning, 2) distributed edge model training at the edge, and 3) edge model inference. Next, we identify general design principles of green edge AI according to the main energy consumption components of practical edge AI systems. These principles provide clear threads to discuss and classify the existing design methodologies for achieving energy-efficient edge AI. In addition, we elaborate on several potential research directions and opportunities that could further improve the EE of edge AI. It is worth noting that a closely related survey on green and energy-efficient AI was presented in [43]. However, it focused on cloud AI and did not address the distinctive challenges in MEC environments. Also, while the discussions in [36, 37, 44, 45] cover all the three key tasks of edge AI to certain extents, the EE-oriented design approaches fall outsite their primary scopes.
@@ I-D Paper Organization
The rest of this paper is organized as follows. In Section II, we first introduce the three key tasks of edge AI, followed by a summary of the general design principles of green edge AI. In Sections III, IV, and V, we respectively review the EE-oriented design approaches for the three key tasks of edge AI, namely data acquisition for centralized edge learning, distributed edge model training, and edge model inference, respectively. Section VI highlights the potential future research directions, and Section VII concludes this paper.
Figure 1: Illustration of the three key tasks in edge AI, including: (Left) data acquisition for centralized edge learning, (Middle) distributed edge model training, (Right) edge model inference.
@@ II Key Tasks of Edge AI and General Design Principles of Green Edge AI
In this section, we first introduce the three key tasks in edge AI and analyze their respective sources of energy consumption. We then decompose the energy consumption of edge AI systems into sensing, communication, and computation energy consumption, based on the configurations of realistic sensing modules, communication technologies, and edge AI platforms. From this analysis, we identify general design principles for green edge AI.
@@ II-A Key Tasks of Edge AI
Edge AI Task
Sensing
Communication
Computation
Data Acquistion for Centralized Edge Model Training
Distributed Edge Model Training
Edge Model Inference
@@ II-B Breakdown of Energy Consumption
Sensing Energy: In TABLE III, the sensing power consumption of four sensing modules is presented, including a 5MP imaging sensor [52], a LiDAR camera [53], a radar sensor [54], and an optical sensor for heart rate monitoring [55]. It can be conclued from TABLE III that the energy consumed during sensing operations is primarily influenced by the sensing mechanism, sampling rate, and sampling quality. For instance, an imaging sensor can consume much higher power (96 mW) compared to an optical sensor for heart rate monitoring (0.72 mW). Besides, the power consumption of a sensing module in active and standby modes are in a sharp difference, e.g., 96 mW and 165 $\mu$W respectively for the imaging sensor, implying the significance of sampling frequency. Furthermore, configurating the LiDAR camera with different depth resolutions (i.e., VGA and XGA) results in varying power consumption. This fact is also exemplified by the case of the radar sensor, again confirming the critical impact of sampling quality on sensing energy consumption.
@@
Communication Energy: Efficient and reliable communication between MEC servers and EUDs is crucial for their collaborative functioning in all three key tasks of edge AI. The power consumption associated with communication depends on many factors, such as carrier frequency, modulation scheme, transmission data rate, and antenna deployment. For instance, an EUD can consume up to 15 dBm ($\sim$32 mW) in a 2.4 GHz WiFi network [56]. More recently, fifth-generation (5G) new radio (NR) offers two frequency modes in sub-6 GHz and millimeter-wave bands, allowing EUDs transmit at the maximum power of 23 dBm ($\sim$199 mW) and 26 dBm ($\sim$398 mW) respectively [57]. While numerous communication technologies have been developed, Bluetooth, WiFi, and cellular communications are particularly suitable for edge AI applications, considering the requirements for data rate and transmission range. To achieve energy-efficient edge AI, it is crucial to select the appropriate communication technology that meets the application’s needs while minimizing power consumption. Moreover, rather than perceiving the communication channel as a mere conduit for data transmission, it is necessary to adapt existing communication protocols to the specific characteristics of edge AI tasks.
@@
Computation Energy: The majority of computations performed by EUDs in edge AI systems are dedicated to distributed model training and model inference. As depicted in TABLE IV, the power consumption associated with computation varies across different edge AI platforms, typically ranging from a few to a few tens of Watts. While the exact power consumption is affected by numerous factors, there is generally a positive correlation between processing speed and power consumption. Additionally, the types of AI workloads directly impact the power consumption. For instance, training a ResNet-110 [58] and VGG-16 [59] model on the NVIDIA Jeston TX2 platform requires approximately $8\times 10^{5}$ and $3.8\times 10^{5}$ Joules of energy, respectively [60]. These energy requirements correspond to approximately 13 and 6.5 times the battery capacity of an iPhone 14. Similarly, utilizing different DNN models for inference will result in varying levels of computation energy consumption. Furthermore, it has been observed that specialized AI accelerators, such as tensor processing units (TPUs), are more energy-efficient compared to general-purpose graphics processing units (GPUs).
@@
In conclusion, the energy consumption in edge AI systems is influenced by different factors for each of the three sources discussed. Consequently, in order to mitigate the energy consumption associated with sensing, communication, and computation, customized approaches need to be developed based on the overarching design principles for energy-efficient edge AI, which will be explored in the next subsection.
@@
TABLE IV: Computational Resources, Processing Speed, and Computation Power Consumption of Different Edge AI ${}^{\dagger}$
Model
RPi-4B
Jetson Nano
Jetson TX2
Coral Dev Board
CPU
Broadcom BCM2711 (Quad core Cortex-A72)
Quad-core ARM Cortex-A57
Dual-Core NVIDIA Denver 2, Quad-core ARM Cortex-A57
NXP i.MX 8M SoC (Quad Cortex-A53, Cortex-M4F)
GPU
N.A.
128 NVIDIA CUDA cores
256 NVIDIA CUDA cores
Integrated GC7000 Lite Graphics
TPU
N.A.
N.A.
N.A.
Google Edge TPU
Memory
1/2/4/8GB LPDDR4-3200 SDRAM
4 GB 64-bit LPDDR4
8GB 128-bit LPDDR4 Memory
1/4 GB LPDDR4
Processing speed
9.9 GFLOPS
472 GFLOPS
1.33 TFLOPS
4 TFOPS
Power consumption
$\sim$7.3 W
5$\sim$10 W
7.5$\sim$15 W
4 W

### v2

@@ Abstract
Artificial intelligence (AI) technologies have emerged as pivotal enablers across a multitude of industries, including consumer electronics, healthcare, and manufacturing, largely due to their significant resurgence over the past decade. The transformative power of AI is primarily derived from the utilization of deep neural networks (DNNs), which require extensive data for training and substantial computational resources for processing. Consequently, DNN models are typically trained and deployed on resource-rich cloud servers. However, due to potential latency issues associated with cloud communications, deep learning (DL) workflows (e.g., DNN training and inference) are increasingly being transitioned to wireless edge networks in proximity to end-user devices (EUDs). This shift is designed to support latency-sensitive applications and has given rise to a new paradigm of edge AI, which will play a critical role in upcoming sixth-generation (6G) networks to support ubiquitous AI applications. Despite its considerable potential, edge AI faces substantial challenges, mostly due to the dichotomy between the resource limitations of wireless edge networks and the resource-intensive nature of DL. Specifically, the acquisition of large-scale data, as well as the training and inference processes of DNNs, can rapidly deplete the battery energy of EUDs. This necessitates an energy-conscious approach to edge AI to ensure both optimal and sustainable performance. In this paper, we present a contemporary survey on green edge AI. We commence by analyzing the principal energy consumption components of edge AI systems to identify the fundamental design principles of green edge AI. Guided by these principles, we then explore energy-efficient design methodologies for the three critical tasks in edge AI systems, including training data acquisition, edge training, and edge inference. Finally, we underscore potential future research directions to further enhance the energy efficiency of edge AI.
@@ I-A Background
Six decades after its inception at the 1956 Dartmouth Conference [1], artificial intelligence (AI) demonstrated its potential to surpass human intelligence in 2016. This was evidenced when a computer program named AlphaGo [2] triumphed over a top-ranked human Go player for the first time, which not only signified a landmark moment in AI history but also catalyzed a resurgence in AI interest and development. The AI renaissance has also been fueled by advancements in computing hardware, the proliferation of big data, and the advent of deep neural networks (DNNs). As a result, AI technologies have made remarkable breakthroughs across a multitude of disciplines, including but not limited to, computer vision [3], natural language processing (NLP) [4], healthcare [5], manufacturing [6], and financial technology (FinTech) [7]. Consequently, AI is rapidly emerging as a restless engine of future productivity and has been recognized as a national strategic priority by major global economies, including China, the United States, and the European Union.
@@
The success of AI is underpinned by the power of advanced data analytics, which typically entails substantial computation overhead [8]. Traditionally, AI computation workloads (e.g., DNN training and inference) initiated by end-user devices (EUDs) are offloaded to the Cloud, which boasts virtually unlimited computational resources [9]. However, the prevalent cloud AI paradigm falls short in supporting latency-critical applications. This is due to the multi-hop routing and network congestion between EUDs and remote public cloud servers, which can result in round-trip time (RTT) of tens to hundreds of milliseconds [10]. Furthermore, a cloud server can become inundated by a high volume of concurrent AI service requests, leading to scalability issues [11]. Concentrating all computational resources in a few cloud data centers also exposes the system to single-point failures and security threats. Moreover, data collected from EUDs may be at risk of privacy breaches [12]. These significant limitations of cloud AI have necessitated the exploration of alternative solutions for delivering AI services.
@@
Gratefully, the drawbacks of cloud AI can be complemented by the emerging mobile edge computing (MEC) technologies, which integrate cloud-like functionalities in the radio access network (RAN) near EUDs [13]. The synergy between MEC and AI enables a major paradigm shift, from cloud AI toward edge AI [14, 15], where the AI computations of EUDs can be offloaded to a wirelessly connected MEC server. As such, the data communication latency can be trimmed significantly to the millisecond level [16], which is essential for real-time interactive AI applications such as mixed reality, smart robots, and autonomous vehicles. Edge AI was identified by Huawei, a leading provider of telecom infrastructures and devices, as one of the six pillars in the sixth-generation (6G) wireless networks [11]. The International Telecommunication Union Radiocommunication Sector (ITU-R) has also defined integrated AI and communication as a key usage scenario of IMT-2030 (6G) in June 2023 [17]. Since 6G is expected to rollout by 2030, edge AI technologies are foreseen to receive unprecedented attention in the near future as predicted by the Gartner Hype Cycle [18]. Apart from latency reduction, edge AI also enjoys many other benefits over cloud AI [19, 20]. On the one hand, the relatively low investment of MEC infrastructures allows their deployments at scale so that the massive AI service demands in 6G can be accommodated efficiently. On the other hand, edge AI inherits the privacy and security enhancements in MEC, since the computing infrastructures are highly distributed and can be privately owned. For instance, without relying on the cloud service providers, user data can be kept and processed inside the local domains, such as university and enterprise networks, reducing the chance of information leakage. Nevertheless, despite all these merits, there exists a significant energy challenge of edge AI, which is caused by the rapidly growing power consumption of AI workloads and the limited energy resources over the wireless edge network [21].
@@ I-B The Energy Challenge in Edge AI
As complex DL processing tasks move away from the Cloud [37], edge AI is emerging as a more sustainable solution for intelligent mobile applications. To a great extent, this is due to the use of low-power edge computing devices, which saves the high energy costs associated with operating cloud data centers. However, the energy challenge remains a significant issue in edge AI, and in fact, becomes even more pronounced [19]. Firstly, the number of EUDs requiring AI services, such as smartphones, personal computers, and vehicles, is increasing at an astonishing rate. A market analysis report by Tractica predicted that the annual worldwide shipment of AI-enabled EUDs will reach 2.6 billion units by 2025, representing a 16-fold increase compared to 2018 [38]. The widespread adoption of intelligent mobile applications will undoubtedly exacerbate the energy crisis of edge AI. Secondly, EUDs are primarily powered by batteries, the development of which significantly lags behind AI algorithmic innovations. Consequently, if compute-intensive DL algorithms are executed on EUDs, they can quickly deplete the battery energy. This not only hinders the provision of stable AI services but may also disrupt other regular system operations. While such an issue can be mitigated with MEC infrastructures, it is challenging to provide sufficient resources to meet the heavy AI computation workloads offloaded from a large number of EUDs. Thirdly, the collaboration between EUDs and MEC servers involves frequent two-way wireless communications [39]. As a result, the energy consumption of communication contributes significantly to the overall energy consumption of edge AI systems, in addition to the computation energy consumption in the mobile AI paradigm, where EUDs handle all the computation workloads.
Figure 1: Paper organization.
To provide high-quality, low-latency, and sustainable AI services to EUDs, it is crucial to develop dedicated solutions to address the energy challenge in edge AI. This forms the primary motivation for this survey on green edge AI. As the first survey of its kind, our focus will be on reducing energy consumption at EUDs, given that MEC infrastructures are likely to have stable power supplies. The improvements of EE shall naturally relieve the environmental impacts of edge AI due to the reduced demand of conventional non-renewable energy.
@@
Ref. AI Tasks Energy Efficiency- oriented Device- edge Cooperation Overview
Dataset Acqui. Model Training DNN Inference
@@
[47] $\bm{\times}$$\bm{\times}$ $\bm{\checkmark}$$\bm{\times}$ $\bm{\checkmark}$ Emerging trends, tools, and techniques for efficient DNN inference on EUDs, including designs of architectures, algorithms, hardware, and accelerators
[48] $\bm{\checkmark}$ $\bm{\checkmark}$ $\bm{\checkmark}$ $\bm{\checkmark}$$\bm{\times}$ A systematic overview on green deep learning techniques, including energy- efficient training, energy-efficient inference, and efficient data usage strategies
@@
Ours $\bm{\checkmark}$ $\bm{\checkmark}$ $\bm{\checkmark}$ $\bm{\checkmark}$ $\bm{\checkmark}$ Comprehensive literature review on EE-oriented design approaches for edge AI, including training data acquisition, distributed edge training, and edge inference
I-C Related Survey Papers and Our Contributions
There are several recent survey papers on edge AI [14, 15, 16, 39, 40, 41, 42]. Specifically, a survey on edge AI was first presented in [14], which provides an overview of the key architectures and technologies for training and inference at the mobile edge network. Applications and open challenges of the synergy between DL and MEC were discussed in [16]. The communication challenges of edge AI were analyzed in [39], for which, numerous communication-efficient techniques were introduced. Similarly, practical edge AI techniques and their interplays with advanced wireless communication systems were elaborated in [40]. The challenges of edge AI were also examined from the general perspective of resource scarcity in [41]. Scalable and trustworthy edge AI systems, built upon the latest innovations in 6G and distributed machine learning (ML) technologies, were discussed in [15]. Additionally, the authors of [42] identified the fundamental components of edge AI as edge caching, edge training, edge inference, and edge offloading, and summarized the respective research results.
@@ I-C Related Survey Papers and Our Contributions
This paper presents a timely and comprehensive literature review on EE-oriented design approaches for edge AI. The novel contributions of this study, compared with existing survey papers on edge AI, are summarized in Table I. Specifically, we first introduce the three key tasks of edge AI, namely 1) data acquisition for centralized edge learning, 2) distributed edge model training, and 3) edge model inference. Next, we identify general design principles of green edge AI according to the main energy consumption components of practical edge AI systems. These principles provide clear threads to discuss and classify the existing design methodologies for achieving energy-efficient edge AI. In addition, we elaborate on several potential research directions and opportunities that could further improve the EE of edge AI. It is worth noting that a closely related survey on green and energy-efficient AI was presented in [48]. However, it focused on cloud AI and did not address the distinctive challenges in MEC environments. Also, while the discussions in [41, 42, 49, 50] cover all three key tasks of edge AI to certain extents, the EE-oriented design approaches fall outside their primary scopes.
@@ I-D Paper Organization
The rest of this paper is organized as follows. In Section II, we first introduce the three key tasks of edge AI, followed by a summary of the general design principles of green edge AI. In Sections III, IV, and V, we respectively review the EE-oriented design approaches for the three edge AI tasks, namely data acquisition for centralized edge learning, distributed edge model training, and edge model inference. Section VI highlights the potential future research directions of green edge AI, and Section VII concludes this paper. Fig. 1 summarizes the paper organization.
@@ II Key Tasks of Edge AI and General Design Principles of Green Edge AI
In this section, we first introduce the three key tasks in edge AI systems and analyze their respective sources of energy consumption. We then decompose the energy consumption of edge AI systems into sensing [^2] Sensing refers to the process of data sampling on EUDs in this paper., communication, and computation energy consumption, based on the configurations of realistic sensing modules, communication technologies, and edge AI platforms. From this analysis, we identify the general design principles for green edge AI.
@@ II-A Key Tasks of Edge AI
Figure 2: Illustration of the three key tasks in edge AI systems, including: (Left) data acquisition for centralized edge learning, (Middle) distributed edge model training, (Right) edge model inference.
@@
Edge AI Task Sensing Communication Computation
Data Acquistion for Centralized Edge Model Training $\bm{\checkmark}$ $\bm{\checkmark}$$\bm{\times}$
Distributed Edge Model Training $\bm{\times}$ $\bm{\checkmark}$ $\bm{\checkmark}$
Edge Model Inference $\bm{\times}$ $\bm{\checkmark}$ $\bm{\checkmark}$
@@ II-B Breakdown of Energy Consumption
Sensing Energy: In TABLE III, the sensing power consumption of four sensing modules is presented, including a 5 megapixel (MP) imaging sensor [57], a LiDAR camera [58], a radar sensor [59], and an optical sensor for heart rate monitoring [60]. It can be concluded from TABLE III that the energy consumed during sensing operations is primarily influenced by the sensing mechanism, sampling rate, and sampling quality. For instance, an imaging sensor can consume much higher power (96 mW) compared to an optical sensor for heart rate monitoring (0.72 mW). Besides, the power consumption of a sensing module in active and standby modes are in a sharp difference, e.g., 96 mW and 165 $\mu$W respectively for the imaging sensor, implying the significance of sampling frequency. Furthermore, configurating the LiDAR camera with different depth resolutions (i.e., VGA and XGA) results in varying power consumption. This fact is also exemplified by the case of the radar sensor, again confirming the critical impact of sampling quality on sensing energy consumption.
@@
Communication Energy: Efficient and reliable communication between MEC servers and EUDs is crucial for their collaborative functioning in all three key tasks of edge AI. The power consumption associated with communication depends on many factors, such as carrier frequency, modulation scheme, transmission data rate, and antenna deployment. For instance, an EUD can consume up to 15 dBm ($\sim$32 mW) in a 2.4 GHz WiFi network [61]. More recently, fifth-generation (5G) new radio (NR) offers two frequency modes in sub-6 GHz and millimeter-wave bands, allowing EUDs transmit at the maximum power of 23 dBm ($\sim$199 mW) and 26 dBm ($\sim$398 mW) respectively [62]. While numerous communication technologies have been developed, Bluetooth, WiFi, and cellular communication are particularly suitable for edge AI applications, considering the requirements for data rate and transmission range. To achieve energy-efficient edge AI, it is crucial to select the appropriate communication technology that meets the application’s needs while minimizing power consumption. Moreover, rather than perceiving the communication channel as a mere conduit for data transmission, it is necessary to adapt existing communication protocols to the specific characteristics of edge AI tasks.
@@
Computation Energy: The majority of computations performed by EUDs in edge AI systems are dedicated to distributed model training and model inference. As depicted in TABLE IV, the power consumption associated with computation varies across different edge AI platforms, typically ranging from a few to a few tens of Watts. While the exact power consumption is affected by numerous factors, there is generally a positive correlation between processing speed and power consumption. Additionally, the types of AI workloads directly impact the power consumption. For instance, training a ResNet-110 [63] and VGG-16 [64] model on the NVIDIA Jeston TX2 platform requires approximately $8\times 10^{5}$ and $3.8\times 10^{5}$ Joules of energy, respectively [65]. These energy requirements correspond to approximately 17.5 and 8.3 times the battery capacity of an iPhone 14 (i.e., 3279 mAh operating at 3.87 V). Similarly, utilizing different DNN models for inference results in varying levels of computation energy consumption. Furthermore, it has been observed that specialized AI accelerators, such as tensor processing units (TPUs), are more energy-efficient compared to general-purpose graphics processing units (GPUs).
@@
In conclusion, the energy consumption in edge AI systems is influenced by different factors for each of the three sources discussed. Moreover, for each edge AI task, there is a distinctive interplay among different components of energy consumption as will be detailed in the sequel. Consequently, aiming at minimizing energy consumption associated with sensing, communication, and computation, customized approaches need to be developed based on the overarching design principles for energy-efficient edge AI, which will be explored in the next subsection.
@@
TABLE IV: Computational Resources, Processing Speed, and Computation Power Consumption of Different Edge AI Platforms
Model RPi-4B Jetson Nano Jetson TX2 Coral Dev Board
CPU Broadcom BCM2711 (Quad-core Cortex-A72) Quad-core ARM Cortex-A57 MPCore processor Dual-core NVIDIA Denver 2 64-bit CPU and Quad-core ARM Cortex-A57 MPCore processor NXP i.MX 8M SoC (Quad Cortex-A53, Cortex-M4F)
GPU N.A. NVIDIA Maxwell architecture GPU with 128 CUDA cores NVIDIA Pascal architecture GPU with 256 CUDA cores Integrated GC7000 Lite Graphics
TPU N.A. N.A. N.A. Google Edge TPU
Memory 1/2/4/8GB LPDDR4 -3200 SDRAM 4 GB 64-bit LPDDR4 8GB 128-bit LPDDR4 1/2/4 GB LPDDR4
Processing speed 9.69 GFLOPS 472 GFLOPS 1.33 TFLOPS 4 TFOPS
Power consumption Idle: $2.8$ W Average: $7.2$ W Maximum: $8.2$ W 5$\sim$10 W 7.5$\sim$15 W 4 W
  • The processing speed and power consumption of RPi-4B are measurement results of the 4 GB memory model available at: https://web.eece.maine.edu/~vweaver/group/green_machines.html. The processing speed and power consumption of Jeston Nano and Jetson TX2 are available at the official website of NVIDIA: https://developer.nvidia.com/embedded/jetson-modules. The processing speed and power consumption of the Cora Dev Board are based on the specification of the Google Edge TPU available at: https://coral.ai/static/files/Coral-Dev-Board-datasheet.pdf.